I am a 2-digit number. The sum of my digits is 11. The tens digit is odd. The ones digit is 3 less than the tens digit. What number am I?

Answer by stanbon(75887)

I am a 2-digit number. The sum of my digits is 11. The tens digit is odd. The ones digit is 3 less than the tens digit. What number am I?
--------------------
Let the number be 10t+u
--------------------
Equations:
t + u = 11
u = t -3
--------------
Substitute for "u" and solve for "t":
t + t-3 = 11
2t = 14
t = 7
---
Since t+u = 11, u = 4
---------------------------
The number is 74
=================================
Cheers,
Stan H.